OpenStudy (anonymous):

What is the equation of the line that passes through the point (3, 2) and has a slope of 4? Answer y = 4x - 10 y = 4x + 1 y = 4x - 1 y = 4x - 5

OpenStudy (anonymous):

y = 4x - 10

OpenStudy (karatechopper):

y-2=4(x-3) y-2=4x-12 y-2+2=4x-12+2 y=4x-10 ur answer is y=4x-10